OpenStudy (anonymous):

Really need help Evaluate the function f(x) = x 2 + 1 for f(2). a.1 b.5 c.7

OpenStudy (snuggielad):

In an F of X problem all you do is evaluate the problem replacing the x with the number inside of the parenthesis. So, for this it would be f(2)=2(2) + 1

OpenStudy (anonymous):

so how do i get the answer?

OpenStudy (snuggielad):

Work that out...Forget the left side work out the right side...you know what x is but now you need to know the combination of F of X

OpenStudy (snuggielad):

So multiply two and two then add one

OpenStudy (snuggielad):

It is easy and sometimes fun when you get the hang of it. You are in algebra one I can tell but when you get into algebra two you will start graphing them which is much more difficult.

OpenStudy (snuggielad):

Did you get it?

OpenStudy (snuggielad):

It is like this...(blank)=a+1 if a = 1...the answer would be found by replacing the in the right side of the problem with 1

OpenStudy (snuggielad):

If you are waiting on me to give you an answer you won't get it I am sorry.

OpenStudy (anonymous):

sorry i was away for a bit give me a sec

OpenStudy (anonymous):

so i got 5 as the answer. is that correct?